Regular bananas sell for $0.49 per pound. Organic bananas sell for $0.89 per pound. If Kate plans to purchase 2 1/2 pounds and sales tax 5%, how much would she pay for the organic bananas, including sales tax?
Answer by ikleyn(52786)

.
Regular bananas sell for $0.49 per pound. Organic bananas sell for $0.89 per pound.
If Kate plans to purchase 2 1/2 pounds and sales tax 5%, how much would she pay for the organic bananas,
including sales tax?
~~~~~~~~~~~~~~~~~~


2 1/2 pounds = 2.5 pounds.


She should pay  2.5*0.89 + 0.05*2.5*0.89 = 2.5*0.89*(1+0.05) = 2.5*0.89*1.05 = 2.34 dollars for her purchase, 

                             including the tax.    ANSWER

Solved and carefully explained.